OpenStudy (anonymous):

solve (13+3x)^1/2-5=0

OpenStudy (anonymous):

-4, -4 and 4, -3, -3 and -4

OpenStudy (anonymous):

last one for me. \[(13+3x)=5^2=25\] \[3x=12\] \[x=4\] good luck with the rest.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

gnight

OpenStudy (anonymous):

night